Find the slope of the line passing through the points (–4, 1) and (–4, 4).
Answer by sdmmadam@yahoo.com(530)

Find the slope of the line passing through the points (–4, 1) and (–4, 4).
Observe that the two given points are
points on the vertical line
x= -4
(with one point 1 unit above the x-axis
and the other point 4 units above the x-axis on this line)
Any line vertical (parallell to the y-axis) is perpendicular to the x-axis and hence makes 90 degrees with the x-axis and tan(90) is infinity and hence slope of the given line is infinity.
Note; If you wish to apply the formula for slope of the line joining two given points P(x1,y1) and Q(x2,y2) given by
slope = (y2-y1)/(x2-x1)
here the given points being P(–4, 1) and Q(–4, 4)(with x1=-4,y1=1;x2=-4,y2=4
and therefore slope =(4-1)/[(-4)-(-4)]= 3/(-4+4)=3/0 = infinity
Note: Please draw the line and understand. It is much easier that way. I am sorry I don't have the facility to draw the diagram in the answer box!